Using power series method,solve the following differential equation

(x3+3)y"+4xy'+y=0


1
Expert's answer
2022-02-02T15:13:54-0500

Let "y = \sum_{n=0}^{\infty}a_nx^n"

Then "y' = \sum_{n=1}^{\infty}na_nx^{n-1}"

And "y'' = \sum_{n=2}^{\infty}n(n-1)a_nx^{n-2}"


Then we have that

"(x^3+3)y'' = \sum_{n=2}^{\infty}n(n-1)a_nx^{n+1} + \sum_{n=2}^{\infty}3n(n-1)a_nx^{n-2} ... (*)"


Also

"4xy' = \sum_{n=1}^{\infty}4na_nx^{n} ... (**)"


Adding up and simplifying y, (*) and (**), we have

"\sum_{n=3}^{\infty}(n-1)(n-2)a_{n-1}x^{n} + \sum_{n=0}^{\infty}3(n+1)(n+2)a_{n+2}x^{n}"

+"a_0 + \sum_{n=1}^{\infty}(4n+1)a_nx^n = 0"


Simplifying further, we have

"\sum_{n=3}^{\infty} \big[(n-1)(n-2)a_{n-1} +3(n+1)(n+2)a_{n+2} +(4n+1)a_n \big]x_n"

"+ 6a_2 + 18a_3x+36a_4x^2+a_0+5a_1x+9a_2x^2 = 0"


Considering

"\sum_{n=3}^{\infty} \big[(n-1)(n-2)a_{n-1} +3(n+1)(n+2)a_{n+2} +(4n+1)a_n \big]x_n"

when "n \geq 1," we have that


"a_{n+2} = \dfrac{(1-n)(n-2)a_{n-1} - (4n+1)a_n}{3(n+1)(n+2)}"


So the solution to the differential equation is

"y = a_0 + a_1x +a_2x^2" "+ \sum_{n=1}^{\infty} \big[ \dfrac{(1-n)(n-2)a_{n-1} - (4n+1)a_n}{3(n+1)(n+2)} \big]x^{n+2}"